Please help me with the following problem:
Joe’s Used Cars sold 50 cars in 1999. In 2004 they sold 180 cars.

Find the average rate of change in cars sold from 1999 to 2004.
Thanks

Answer by gonzo(654)

in 1999 he sold 50
in 2004 he sold 180
change in number of cars sold is 180 - 50 = 130.
change in number of years is 2004 - 1999 = 5
change in number of cars sold per year is 130 / 5 = 26